# Vector Quantized-Elites: Unsupervised and Problem-Agnostic Quality-Diversity Optimization

### v1

@@ Abstract
Quality-Diversity algorithms have transformed optimization by prioritizing the discovery of diverse, high-performing solutions over a single optimal result. However, traditional Quality-Diversity methods, such as MAP-Elites, rely heavily on predefined behavioral descriptors and complete prior knowledge of the task to define the behavioral space grid, limiting their flexibility and applicability. In this work, we introduce Vector Quantized-Elites (VQ-Elites), a novel Quality-Diversity algorithm that autonomously constructs a structured behavioral space grid using unsupervised learning, eliminating the need for prior task-specific knowledge. At the core of VQ-Elites is the integration of Vector Quantized Variational Autoencoders, which enables the dynamic learning of behavioral descriptors and the generation of a structured, rather than unstructured, behavioral space grid — a significant advancement over existing unsupervised Quality-Diversity approaches. This design establishes VQ-Elites as a flexible, robust, and task-agnostic optimization framework. To further enhance the performance of unsupervised Quality-Diversity algorithms, we introduce two key components: behavioral space bounding and cooperation mechanisms, which significantly improve convergence and performance. We validate VQ-Elites on robotic arm pose-reaching and mobile robot space-covering tasks. The results demonstrate its ability to efficiently generate diverse, high-quality solutions, emphasizing its adaptability, scalability, robustness to hyperparameters, and potential to extend Quality-Diversity optimization to complex, previously inaccessible domains.
@@ I Introduction
Figure 1: Overview of VQ-Elites
@@
Quality-Diversity (QD) algorithms address the limitations of traditional evolutionary algorithms by redefining the optimization objective [1, 2]. Rather than seeking a single optimal solution, QD algorithms, such as MAP-Elites [3, 4], aim to uncover a diverse set of high-performing solutions across a broad behavioral landscape. This “illumination” approach generates multiple effective parameter vectors, each associated with a unique behavior. By promoting both diversity and quality, QD algorithms enable more flexible and robust applications, particularly in autonomous systems, where a repertoire of high-quality behaviors is essential for adapting to dynamic challenges. For instance, QD algorithms have demonstrated the ability to generate diverse locomotion behaviors in quadruped robots [4, 5, 6].
Despite these advances, QD algorithms face significant limitations, particularly regarding the reliance on predefined or ”hard-coded” behavioral descriptors (BDs). In MAP-Elites, for example, the behavior space is often discretized using a set of user-defined descriptors to categorize behaviors, which imposes constraints on the algorithm’s adaptability. These hard-coded descriptors assume a priori knowledge about the task or environment and may not fully capture the complexity or diversity inherent in high-dimensional, novel tasks. This constraint limits the algorithm’s flexibility, as predefined descriptors may inadvertently filter out behaviors that fall outside of the original descriptors or misrepresent complex solutions. Moreover, for tasks where the behavior descriptors are hard to design or even unknown, the manual design of BDs becomes impractical and can hinder the QD algorithm’s capacity to scale or generalize to new domains.
Early approaches to automatic BD extraction relied on large static datasets to learn lower-dimensional representations using techniques like Principal Component Analysis (PCA) or auto-encoders. For instance, Cully and Demiris [7] utilized a dataset of handwritten digits to train an auto-encoder, creating a low-dimensional latent space that captured the key features of different digits. This latent space then served as the behavioral descriptor space, enabling QD algorithms to generate a diverse set of solutions spanning the learned feature space. This approach allowed a robotic arm to autonomously learn to draw digits without requiring manual definitions of their key characteristics.
@@
To demonstrate the effectiveness of VQ-Elites, we evaluate its performance across a variety of tasks that highlight its adaptability and robustness in complex scenarios. These include both structured and unstructured space-covering tasks, as well as high-dimensional pose-reaching tasks for robotic manipulators. In each experiment, we assess VQ-Elites’ ability to generate diverse, high-quality behaviors while maintaining an organized archive. The results underscore its strengths in flexibility and robustness, showcasing its potential to advance QD frameworks for autonomous robotic systems and extend their applicability to more complex problem domains.
@@
  • We introduce two components, Behavioral Space Bounding and Cooperation, which drastically improve the convergence of unsupervised QD algorithms.
  • We evaluate VQ-Elites across diverse tasks, including space-covering challenges, pose-reaching tasks for robotic manipulators, and scenarios involving high-dimensional data, and constrained systems.
  • Experimental results show that VQ-Elites efficiently generates diverse, high-quality behaviors while preserving a well-organized archive, highlighting its potential to enhance QD frameworks for autonomous robotic systems and extend their applicability to complex problem domains.
@@ II-A Quality-Diversity Algorithms
One of the notable strengths of QD algorithms is their ability to provide multiple high-performing solutions, offering users significant flexibility. This flexibility arises from the wide range of solutions available, allowing users to balance complexity, desired behavior, and practical implementation.
To express this more formally, consider an objective function that returns a fitness value, $f_{\theta}$, and a behavioral descriptor vector, $\boldsymbol{b}_{\theta}$, which defines the behavior:
@@
In practice, the behavioral descriptor characterizes the solution, while the fitness value measures its quality. To clarify the concept of a BD and the difference to the fitness, consider a scenario where a robot must traverse various locations within an environment. In this case, the robot’s coordinates can be the BD, while the average speed it reaches the targets the fitness value. Bringing these components together, the optimization algorithm generates multiple solutions, each defined by its BD vector, which describes the behavior or outcome, and its fitness value, which quantifies its effectiveness.
But, what is our ultimate goal when using a Quality-Diversity algorithm? Let us assume, without the loss of generality, that we have a behavior space (feature space) $\mathcal{B}$ and a fitness function that we want to maximize. The ultimate goal of QD optimization is to find all parameters, $\boldsymbol{\theta}$, that maximize the fitness function for every $\boldsymbol{b}\in\mathcal{B}$:
@@
Archives in QD algorithms can take various forms. Perphaps the easiest one involves discretizing the behavior space into a grid, where each cell corresponds to a unique solution. Solutions generated by the algorithm are assigned to the closest cell based on the Euclidean distance between the solution’s BD and the cell’s center. This grid-based method is widely used, particularly in the MAP-Elites algorithm, one of the most popular QD optimization approaches. Other archive variations include more abstract and unstructured designs. For instance, a user may define a distance threshold, allowing a solution to be stored if its distance to all existing BDs exceeds the threshold, or if it outperforms competing BDs within the threshold.
Algorithm 1 MAP-Elites algorithm
procedure MAP-Elites($[n_{1},...,n_{d}]$)
     $\mathcal{A}\longleftarrow$ create_empty_archive($[n_{1},...,n_{d}]$)
     for $i=1\to G$ do $\triangleright$ Initialization: $G$ random $\boldsymbol{\theta}$
         $\boldsymbol{\theta}=$ random_solution()
         add_to_archive($\boldsymbol{\theta},\mathcal{A}$)
     end for
     for $i=1\to I$ do $\triangleright$ Main loop, $I$ iterations
         $\boldsymbol{\theta}=$ selection($\mathcal{A}$)
         $\boldsymbol{\theta}^{\prime}=$ variation($\boldsymbol{\theta}$)
         add_to_archive($\boldsymbol{\theta}^{\prime},\mathcal{A}$)
     end for
     return $\mathcal{A}$
end procedure
procedure add_to_archive($\boldsymbol{\theta},\mathcal{A}$)
     $(p,\mathbf{b})\longleftarrow$ evaluate($\boldsymbol{\theta}$)
     $c\longleftarrow$ get_cell_index($\mathbf{b}$)
     if $\mathcal{A}(c)=null$ or $\mathcal{A}(c).p<p$ then
         $\mathcal{A}(c)\longleftarrow p,\boldsymbol{\theta}$
     end if
end procedure
As a reference, we provide the pseudo-code of the MAP-Elites algorithm (Algo. 1), one of the first and most used QD algorithms. The popularity of MAP-Elites as a QD algorithm stems from its simplicity, both in terms of understanding and implementation. However, its main challenge lies in discretizing the behavior space, which is highly problem- and domain-dependent. This becomes particularly difficult in scenarios involving high-dimensional features or when the behavior space lacks well-defined constraints.
@@ II-B Unsupervised Representation Learning
Representation learning, is a fundamental concept in machine learning and artificial intelligence, focusing on the development of effective data representations that enhance the performance of machine learning models on various tasks. The goal is to automatically transform raw data, such as images, text, or audio, into a more abstract and useful format that captures the essential features of the data while discarding irrelevant or redundant information. Effective representations can simplify learning tasks, reduce the need for domain-specific feature engineering, and enable generalization across different datasets and tasks. In contrast, traditional approaches relied heavily on manual feature engineering, requiring domain experts to identify and extract useful features. This process, however, is often labor-intensive, prone to biases, and limits scalability across different problem domains. Representation learning automates this process, allowing models to discover and learn optimal features from data autonomously. This paradigm shift has led to breakthroughs in various fields, including natural language processing [10], computer vision [11, 12], and speech recognition [13, 14].
VQ-VAE represents a powerful extension of traditional autoencoders, combining ideas from variational autoencoders (VAEs) and vector quantization. Introduced by Oord et al. [15], VQ-VAE addresses the limitations of continuous latent spaces by discretizing them, allowing the model to learn discrete latent representations that are often more interpretable and suitable for tasks such as generation, compression, and unsupervised learning. VQ-VAE has shown remarkable performance in applications such as image and audio synthesis, video generation, and reinforcement learning [16, 17].
@@
One of the key challenges of minimizing this combined loss is the non-differentiability of the quantization step (i.e., selecting the nearest codebook entry), which prevents gradients from flowing back through the encoder during standard training. To address this, VQ-VAE uses a straight-through estimator for gradient updates. Specifically, during the backward pass, the gradients from the decoder are passed through the quantization operation unchanged, which allows effective training despite the discrete nature of the latent space [15].
@@ III Related Work
The MAP-Elites algorithm, introduced by Mouret and Clune [3], established a new paradigm in evolutionary optimization by illuminating diverse regions of the solution space rather than focusing solely on a single optimal solution. MAP-Elites partitions the behavioral space into user-defined niches and searches for the highest-performing solution within each niche. This approach has shown success in robotic control tasks, constrained optimization, and applications requiring broad behavioral diversity [4]. However, MAP-Elites relies heavily on pre-defined behavioral descriptors (BDs), which limits its flexibility and adaptability to dynamic or complex environments where optimal descriptors are not readily apparent.
@@
More recent work has explored integrating reinforcement learning (RL) techniques with MAP-Elites to improve search efficiency and expand behavioral diversity. For example, methods like PGA-MAP-Elites and QD-PG leverage policy gradients alongside MAP-Elites’ mutation-based search, enabling gradient-informed exploration of high-dimensional behaviors [19, 20]. However, these approaches still rely on predefined descriptors or additional complexity from combining RL methods, which may not be ideal for fully unsupervised BD discovery.
@@
The introduction of unsupervised learning methods into the QD framework, like AURORA [9], has enabled the automatic extracting latent behavior representations, allowing QD algorithms to operate independently of pre-defined descriptors. For instance, AURORA employs neural networks to identify meaningful patterns within the behavior space, yet it lacks a structured archive organization akin to MAP-Elites, which can hinder systematic exploration in complex domains [9] (as we will show in our experiments). The TAXONS algorithm [25], the predecessor of AURORA, also utilizes an autoencoder with RGB images as input to construct a low-dimensional behavior space, or outcome space as they characterize it. The authors, also use two metrics to decide the fitness of a solution, the fitness score and the autoencoder’s reconstruction error, and one of them is chosen for each iteration with a probability of 50%. Similarly, techniques leveraging deep reinforcement learning [26] and self-supervised approaches [27] have demonstrated promise in autonomous skill discovery, but their reliance on task-specific tuning can restrict generalization. Lastly, the Quality-Diversity Meta-Evolution (QD-Meta) framework [28], which evolves a population of QD algorithms to dynamically customize behavior spaces for user-defined meta-objectives. By using the Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES), QD-Meta uses neural networks that transform high-dimensional behavior features into compact descriptors while leveraging reinforcement learning for dynamic parameter control. This approach enables the creation of tailored, diverse, and high-quality solutions across various domains without relying on manually crafted behavior spaces. Our approach, VQ-Elites, builds on these advancements by combining MAP-Elites with a VQ-VAE to autonomously generate structured, and flexible behavior spaces, thereby overcoming the limitations of previous unsupervised methods.
@@ IV Vector-Quantized Elites
The primary goal of this work is to develop a novel QD framework that enables the development of algorithms that are able to operate entirely unsupervised, eliminating the need for prior knowledge about the task, environment, or dynamic constraints. Our proposed VQ-Elites algorithm achieves this by leveraging the agent’s experiences in conjunction with Vector Quantized (VQ) machine learning models. This combination allows VQ-Elites to dynamically and autonomously discover meaningful behavioral descriptors that adapt to the specific characteristics of the problem at hand, ensuring flexibility and scalability across diverse tasks and environments. A detailed overview of the algorithm is provided in Algo. 2 and in Fig. 1.
@@
1:Define: $\mathcal{M},\;\mathcal{C}$triangleright VQ-VAE model, behavioral space grid
2:Initialize: $\mathcal{D}\leftarrow\varnothing$, $\mathcal{M}_{\text{codebook}}$triangleright Initialize an empty dataset, Initialize model’s codebook randomly or using K-Means
@@
4:     $\mathcal{A}\longleftarrow\{\varnothing\}_{|C|}$triangleright Initialize an empty archive of size $|C|$ (number of cells in the grid)
@@
14:     $\mathcal{D}_{\text{exploration}}\leftarrow\pi_{\theta}(\alpha|s)_{K}$triangleright Explore the environment $K$ times using random policies and store the data
15:     $\mathcal{M}$, $\mathcal{D}_{\text{exploration}}$) $\triangleright$ Train the VQ-VAE model using the exploration data
16:     $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$triangleright Set the learned codebook as the initial BD grid
17:     $\mathcal{A}\leftarrow\mathcal{M}_{\text{encoder}}(\mathcal{D}_{\text{exploration}})$triangleright Add random solutions to the archive
@@
23:     $\mathcal{D}^{(i)}\longleftarrow d$triangleright Add the raw data to the training dataset
24:     $\mathbf{bd}\longleftarrow\mathcal{M}_{encoder}(d)$triangleright Extract the behavioral descriptor
@@
27:         $\mathcal{A}(c)\longleftarrow fitness,\boldsymbol{\theta}^{\prime}$triangleright Add solution to the archive
@@
31:     $\mathcal{M}$, $\mathcal{D}$) for $\mathcal{E}$ epochs $\triangleright$ Update the VQ-VAE model
32:     $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$triangleright Set the the updated codebook as the BD grid
@@
A defining feature of the VQ-Elites algorithm is its unsupervised nature, which allows it to autonomously learn behavioral descriptors for the discovered solutions. Unlike traditional QD algorithms, where users must manually define BDs [3, 6] or rely on prior knowledge to generate BDs using clustering methods [18], VQ-Elites removes these limitations. Manual or predefined BD selection can restrict the range of problems the algorithm can address or render it impractical in some cases. To overcome this, VQ-Elites leverages representation learning and unsupervised methods to learn BDs directly from the data.
In this work, we employ VQ-VAEs, a generative model that combines autoencoders with vector quantization to create discrete latent representations. VQ-VAEs learn a codebook of vectors, mapping continuous latent vectors onto discrete codes. For high-dimensional data such as images, videos, or multivariate time-series, this mapping typically occurs along the channel dimension. However, when applied to one-dimensional data, VQ-VAE performs gradient-based clustering of the latent vectors, where the discrete representations correspond to centroids — similar to the K-Means algorithm [29].
@@
Alternative unsupervised methods for BD extraction, such as AURORA [9], face significant challenges when learning from online QD data. The dynamic nature of QD archives means that the behavioral space evolves continuously, making it difficult for these methods to adapt effectively in real-time. Moreover, AURORA relies on additional hyperparameters, such as the minimum distance threshold for storing solutions in its unstructured archive. These hyperparameters can be difficult to tune and may require prior domain knowledge, reducing the method’s generality and scalability. While AURORA can be effective in certain scenarios, its reliance on unstructured archives and extra hyperparameters makes it less suitable for tasks that benefit from a structured, adaptive behavior space like the one provided by MAP-Elites.

### v2

@@ Abstract
Quality-Diversity algorithms have transformed optimization by prioritizing the discovery of diverse, high-performing solutions over a single optimal result. However, traditional Quality-Diversity methods, such as MAP-Elites, rely heavily on predefined behavior descriptors and complete prior knowledge of the task to define the behavior space grid, limiting their flexibility and applicability. In this work, we introduce Vector Quantized-Elites (VQ-Elites), a novel Quality-Diversity algorithm that autonomously constructs a structured behavior space grid using unsupervised learning, eliminating the need for prior task-specific knowledge. At the core of VQ-Elites is the integration of Vector Quantized Variational Autoencoders, which enables the dynamic learning of behavior descriptors and the generation of a structured, rather than unstructured, behavior space grid — a significant advancement over existing unsupervised Quality-Diversity approaches. This design establishes VQ-Elites as a flexible, robust, and task-agnostic optimization framework. To further enhance the performance of unsupervised Quality-Diversity algorithms, we introduce behavior space bounding and cooperation mechanisms, which significantly improve convergence and performance, as well as the Effective Diversity Ratio and Coverage Diversity Score, two novel metrics that quantify the actual diversity in the unsupervised setting. We validate VQ-Elites on robotic arm pose-reaching, mobile robot space-covering, and MiniGrid exploration tasks. The results demonstrate its ability to efficiently generate diverse, high-quality solutions, emphasizing its adaptability, scalability, robustness to hyperparameters, and potential to extend Quality-Diversity optimization to complex, previously inaccessible domains.
@@ I Introduction
Figure 1: Overview of VQ-Elites. The algorithm and each component depicted in the figure are described in detail in Sec. IV.
@@
Quality-Diversity (QD) algorithms address the limitations of traditional evolutionary algorithms by redefining the optimization objective [1, 2]. Rather than seeking a single optimal solution, QD algorithms, such as MAP-Elites [3, 4], aim to uncover a diverse set of high-performing solutions across a broad behavior landscape. This “illumination” approach generates multiple effective individuals, each associated with a unique behavior. By promoting both diversity and quality, QD algorithms enable more flexible and robust applications, particularly in autonomous systems, where a repertoire of high-quality behaviors is essential for adapting to dynamic challenges. For instance, QD algorithms have demonstrated the ability to generate diverse locomotion behaviors in quadruped robots [4, 5, 6].
Despite these advances, QD algorithms face significant limitations, particularly regarding the reliance on predefined or ”hard-coded” behavior descriptors (BDs). In MAP-Elites, for example, the behavior space is often discretized using a set of user-defined descriptors to categorize behaviors, which imposes constraints on the algorithm’s adaptability. These hard-coded descriptors assume a priori knowledge about the task or environment and may not fully capture the complexity or diversity inherent in high-dimensional, novel tasks. This constraint limits the algorithm’s flexibility, as predefined descriptors may inadvertently filter out behaviors that fall outside of the original descriptors or misrepresent complex solutions. Moreover, for tasks where the behavior descriptors are hard to design or even unknown, the manual design of BDs becomes impractical and can hinder the QD algorithm’s capacity to scale or generalize to new domains.
Early approaches to automatic BD extraction relied on large static datasets to learn lower-dimensional representations using techniques like Principal Component Analysis (PCA) or auto-encoders. For instance, Cully and Demiris [7] utilized a dataset of handwritten digits to train an auto-encoder, creating a low-dimensional latent space that captured the key features of different digits. This latent space then served as the behavior descriptor space, enabling QD algorithms to generate a diverse set of solutions spanning the learned feature space. This approach allowed a robotic arm to autonomously learn to draw digits without requiring manual definitions of their key characteristics.
@@
To demonstrate the effectiveness of VQ-Elites, we evaluate its performance across a variety of tasks that highlight its adaptability and robustness in complex scenarios. These include both structured and unstructured space-covering tasks, a hard exploration task, as well as high-dimensional pose-reaching tasks for robotic manipulators. In each experiment, we assess VQ-Elites’ ability to generate diverse, high-quality behaviors while maintaining an organized archive. The results underscore its strengths in flexibility and robustness, showcasing its potential to advance QD frameworks and extend their applicability to more complex problem domains.
@@
  • We introduce two components, Behavior Space Bounding and Cooperation, which drastically improve the convergence of unsupervised QD algorithms.
  • We propose two novel metrics, Effective Diversity Ratio and Coverage Diversity Score, to properly quantify the performance of unsupervised QD approaches.
  • We evaluate VQ-Elites across diverse tasks, including space-covering and hard exploration challenges, pose-reaching tasks for robotic manipulators, and scenarios involving high-dimensional data, and constrained systems.
  • Experimental results show that VQ-Elites efficiently generates diverse, high-quality behaviors while preserving a well-organized archive, highlighting its potential to enhance QD frameworks and extend their applicability to complex problem domains.
@@ II-A Quality-Diversity Algorithms
One of the notable strengths of QD algorithms is their ability to provide multiple high-performing solutions, offering users significant flexibility [1, 2]. This flexibility arises from the wide range of solutions available, allowing users to balance complexity, desired behavior, and practical implementation. Here, we present shortly the QD optimization framework for completeness, but we refer the reader to recent surveys for a more detailed overview [2, 1].
To express this more formally, consider an objective function that returns a fitness value, $f_{\theta}$, and a behavior descriptor vector, $\boldsymbol{b}_{\theta}$, which defines the behavior:
@@
In practice, the behavior descriptor characterizes the solution, while the fitness value measures its quality. To clarify the concept of a BD and the difference to the fitness, consider a scenario where a robot must traverse various locations within an environment. In this case, the robot’s coordinates can be the BD, while the average speed it reaches the targets the fitness value. Bringing these components together, the optimization algorithm generates multiple solutions, each defined by its BD vector, which describes the behavior or outcome, and its fitness value, which quantifies its effectiveness.
But, what is our ultimate goal when using a Quality-Diversity algorithm? Let us assume, without the loss of generality, that we have a behavior space (feature space) $\mathcal{B}$ and a fitness function that we want to maximize. The ultimate goal of QD optimization is to find all individuals [^1] In our experiments, we always use real-valued parameter vectors., $\boldsymbol{\theta}$, that maximize the fitness function for every $\boldsymbol{b}\in\mathcal{B}$:
@@
Archives in QD algorithms can take various forms. Perhaps the easiest one involves discretizing the behavior space into a grid, where each cell corresponds to a unique solution. Solutions generated by the algorithm are assigned to the closest cell based on the Euclidean distance between the solution’s BD and the cell’s center. This grid-based method is widely used, particularly in the MAP-Elites algorithm, one of the most popular QD optimization approaches. Other archive variations include more abstract and unstructured designs. For instance, a user may define a distance threshold, allowing a solution to be stored if its distance to all existing BDs exceeds the threshold, or if it outperforms competing BDs within the threshold.
The MAP-Elites algorithm [3] is one of the first and most popular QD algorithms. The popularity of MAP-Elites stems from its simplicity, both in terms of understanding and implementation. However, its main challenge lies in discretizing the behavior space, which is highly problem- and domain-dependent. This becomes particularly difficult in scenarios involving high-dimensional features or when the behavior space lacks well-defined constraints.
@@ II-B Unsupervised Representation Learning
Representation learning is a fundamental concept in machine learning and artificial intelligence, focusing on the development of effective data representations that enhance the performance of machine learning models on various tasks. The goal is to automatically transform raw data, such as images, text, or audio, into a more abstract and useful format that captures the essential features of the data while discarding irrelevant or redundant information. Effective representations can simplify learning tasks, reduce the need for domain-specific feature engineering, and enable generalization across different datasets and tasks. In contrast, traditional approaches relied heavily on manual feature engineering, requiring domain experts to identify and extract useful features. This process, however, is often labor-intensive, prone to biases, and limits scalability across different problem domains. Representation learning automates this process, allowing models to discover and learn optimal features from data autonomously.
VQ-VAE represents a powerful extension of traditional autoencoders, combining ideas from variational autoencoders (VAEs) and vector quantization. Introduced by Oord et al. [10], VQ-VAE addresses the limitations of continuous latent spaces by discretizing them, allowing the model to learn discrete latent representations that are often more interpretable and suitable for tasks such as generation, compression, and unsupervised learning.
@@
Essentially, despite the reconstruction loss that is present in all the encoder-decoder type of architectures, the two new losses operate in a complementary way. The codebook loss adapts the dictionary, while the commitment loss regularizes the encoder. Thus, the former encourages the codebook embeddings to move closer to the encoder outputs, effectively updating the codebook to represent the data more faithfully. On the other hand, the latter acts in the reverse direction, penalizing the encoder for straying too far from the chosen codebook vectors, therefore ensuring stability and consistency in encoding.
One of the key challenges of minimizing this combined loss is the non-differentiability of the quantization step (i.e., selecting the nearest codebook entry), which prevents gradients from flowing back through the encoder during standard training. To address this, VQ-VAE uses a straight-through estimator for gradient updates. Specifically, during the backward pass, the gradients from the decoder are passed through the quantization operation unchanged, which allows effective training despite the discrete nature of the latent space [10]. This paradigm shift to discrete latent spaces has led to breakthroughs in various fields, including natural language processing [11], computer vision [12, 13], speech recognition [14, 15], and reinforcement learning [16, 17].
@@ III Related Work
The MAP-Elites algorithm, introduced by Mouret and Clune [3], established a new paradigm in evolutionary optimization by illuminating diverse regions of the solution space rather than focusing solely on a single optimal solution. This class of algorithms later became known as Quality-Diversity algorithms [18]. MAP-Elites partitions the behavior space into user-defined niches and searches for the highest-performing solution within each niche. This approach has shown success in robotic control tasks, constrained optimization, and applications requiring broad behavioral diversity [4]. Other notable domains where MAP-Elites has been applied, showcasing its applicability and impact, are regarding the robustness of LLMs in adversarial prompts [19], game generation [20], evolving graph structures [21], and image synthesis [22]. However, MAP-Elites relies heavily on pre-defined behavior descriptors (BDs), which limits its flexibility and adaptability to dynamic or complex environments where optimal descriptors are not readily apparent.
@@
More recent work has explored integrating reinforcement learning (RL) techniques with MAP-Elites to improve search efficiency and expand behavior diversity. For example, methods like PGA-MAP-Elites and QD-PG leverage policy gradients alongside MAP-Elites’ mutation-based search, enabling gradient-informed exploration of high-dimensional behaviors [24, 25]. However, these approaches still rely on predefined descriptors or additional complexity from combining RL methods, which may not be ideal for fully unsupervised BD discovery.
@@
The introduction of unsupervised learning methods into the QD framework, like AURORA [9], has enabled the automatic extracting latent behavior representations, allowing QD algorithms to operate independently of pre-defined descriptors. For instance, AURORA employs neural networks to identify meaningful patterns within the behavior space, yet it lacks a structured archive organization akin to MAP-Elites, which can hinder systematic exploration in complex domains [9] (as we will show in our experiments). The TAXONS algorithm [30], the predecessor of AURORA, also utilizes an autoencoder with RGB images as input to construct a low-dimensional behavior space, or outcome space as they characterize it. The authors, also use two metrics to decide the fitness of a solution, the fitness score and the autoencoder’s reconstruction error, and one of them is chosen for each iteration with a probability of 50%. More recently, STAX [31] operates through an alternating two-phase process: exploration, where it builds a diverse set of policies and learns the behavior representation; and exploitation, where it optimizes the performance of policies that have discovered rewards. Similarly, techniques leveraging deep reinforcement learning [32] and self-supervised approaches [33] have demonstrated promise in autonomous skill discovery, but their reliance on task-specific tuning can restrict generalization. Lastly, the Quality-Diversity Meta-Evolution (QD-Meta) framework [34], which evolves a population of QD algorithms to dynamically customize behavior spaces for user-defined meta-objectives. By using the Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES), QD-Meta uses neural networks that transform high-dimensional behavior features into compact descriptors while leveraging reinforcement learning for dynamic parameter control. This approach enables the creation of tailored, diverse, and high-quality solutions across various domains without relying on manually crafted behavior spaces. Our approach, VQ-Elites, builds on these advancements by combining MAP-Elites with a VQ-VAE to autonomously generate structured, and flexible behavior spaces, thereby overcoming the limitations of previous unsupervised methods.
@@ IV Vector-Quantized Elites
The primary goal of this work is to advance the MAP-Elites to operate entirely unsupervised, eliminating the need for prior knowledge about the task, environment, or dynamic constraints, and in the process to propose a new paradigm for the development of unsupervised QD algorithms. Our proposed VQ-Elites algorithm achieves this by leveraging the agent’s experiences in conjunction with Vector Quantized (VQ) machine learning models. This combination allows VQ-Elites to dynamically and autonomously adapt to the specific characteristics of the problem at hand, ensuring flexibility and scalability across diverse tasks and environments. A detailed overview of the algorithm is provided in Algo. 1 and in Fig. 1.
@@
1:Define: $\mathcal{M},\;\mathcal{C}$ $\triangleright$ VQ-VAE model, behavior space grid
2:Initialize: $\mathcal{D}\leftarrow\varnothing$, $\mathcal{M}_{\text{codebook}}$ $\triangleright$ Initialize an empty dataset, Initialize model’s codebook randomly or using K-Means
@@
4:  $\mathcal{A}\longleftarrow\{\varnothing\}_{|C|}$ $\triangleright$ Initialize an empty archive of size $|C|$ (number of cells in the grid)
@@
14:  $\mathcal{D}_{\text{exploration}}\leftarrow\pi_{\theta}(\alpha|s)_{K}$ $\triangleright$ Explore the environment $K$ times using random policies and store the data
15:  train($\mathcal{M}$, $\mathcal{D}_{\text{exploration}}$) $\triangleright$ Train the VQ-VAE model using the exploration data
16:  $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$ $\triangleright$ Set the learned codebook as the initial BD grid
17:  $\mathcal{A}\leftarrow\mathcal{M}_{\text{encoder}}(\mathcal{D}_{\text{exploration}})$ $\triangleright$ Add random solutions to the archive
@@
23:  $\mathcal{D}^{(i)}\longleftarrow d$ $\triangleright$ Add the raw data to the training dataset
24:  $\mathbf{bd}\longleftarrow\mathcal{M}_{encoder}(d)$ $\triangleright$ Extract the behavior descriptor
@@
27:   $\mathcal{A}(c)\longleftarrow fitness,\boldsymbol{\theta}^{\prime}$ $\triangleright$ Add solution to the archive
@@
31:  train($\mathcal{M}$, $\mathcal{D}$) for $\mathcal{E}$ epochs $\triangleright$ Update the VQ-VAE model
32:  $\mathcal{C}\leftarrow\mathcal{M}_{\text{codebook}}$ $\triangleright$ Set the the updated codebook as the BD grid
@@
A defining feature of the VQ-Elites algorithm is its unsupervised nature, which allows it to autonomously learn behavior descriptors for the discovered solutions. Unlike traditional QD algorithms, where users must manually define BDs [3, 6] or rely on prior knowledge to generate BDs using clustering methods [23], VQ-Elites removes these limitations. Manual or predefined BD selection can restrict the range of problems the algorithm can address or render it impractical in some cases. To overcome this, VQ-Eliteslearns BDs directly from the data.
VQ-VAE is a generative model that combines autoencoders with vector quantization to create discrete latent representations. VQ-VAEs learn a codebook of vectors, mapping continuous latent vectors onto discrete codes. For high-dimensional data such as images, videos, or multivariate time-series, this mapping typically occurs along the channel dimension. However, when applied to one-dimensional data, VQ-VAE performs gradient-based clustering of the latent vectors, where the discrete representations correspond to centroids — similar to the K-Means algorithm [35]. Hence, our algorithm can be used with any data modality, e.g., images, videos, or times-series. However, the latent extraction should be done by a fully connected layer on the flattened features extracted by the previous layers of the encoder.
@@
Alternative unsupervised methods for BD extraction, such as AURORA [9], face significant challenges when learning from online QD data. The dynamic nature of QD archives means that the behavior space evolves continuously, making it difficult for these methods to adapt effectively in real-time. Moreover, AURORA relies on additional hyperparameters, such as the minimum distance threshold for storing solutions in its unstructured archive. These hyperparameters can be difficult to tune and may require prior domain knowledge, reducing the method’s generality and scalability. While AURORA can be effective in certain scenarios, its reliance on unstructured archives and extra hyperparameters makes it less suitable for tasks that benefit from a structured, adaptive behavior space like the one provided by MAP-Elites.